A car is negotiating a curved road of radius r. The road is banked at an angle @.The coefficient of friction between the tyres of the car and the road is u. the maximum safe velocity on ths road is what???

Grade:11

1 Answers

Lennon
6 years ago
The centrifugal force is the force moving outward. The frictional is the force moving inwards to prevent the car from slipping out of the road.Centrifugal force=centripetal force=mv^2/rFrictional force= coefficient of friction x normal force NTherefore,the frictional force must be larger than the centrifugal force to hold it.But,the maximum centrifugal force must be equal to the frictional force. If it exceeds the frictional force it slips.So, we getF(c)max=fmv^2/r=μNmv^2=uNrv^2=uNr/mvmax=√(uNr/m)